OpenStudy (annie96):

The sum of two consecutive odd integers is 80. What is the smaller number? 41 43 37 39 Select the pair of equations whose graphs are perpendicular. 2x-8y=9 12x-3y=7 3x+6y=8 y=2x-8 y=2x-7 x+y=3 2y=-3x+5 2x+3y=4

OpenStudy (ikram002p):

first one u just chose the smallest num O.o the second one find the slops in each case (m1,m2) n the equations that gives u m1×m2=-1 is perpendicular.

OpenStudy (zpupster):

x-1+x-3=80 2x=84 x=42 so 42 -1= 42-3=

OpenStudy (annie96):

39